OpenStudy (anonymous):

solve for P. p+40=800-10P

OpenStudy (anonymous):

i got 76!

OpenStudy (praxer):

p+10p=800-40 11p=760 p=760/11

OpenStudy (praxer):

=69.09

OpenStudy (anonymous):

p + 40 = 800 - 10p 11p + 40 = 800 11p = 760 p = 760/11 = 69.090909...

OpenStudy (anonymous):

Oh I forgot to add 10P+1P, Thanks!

OpenStudy (anonymous):

can i ask other question?

OpenStudy (anonymous):

its similar to this I want to check if my answer is right or not

OpenStudy (praxer):

Surely..:)